# Local analysis of iterative reconstruction from discrete generalized Radon transform data in the plane

### v1

@@ 8.1. Reduction of the DTB function to an integral
(9.5) $\begin{split}&\Delta G_{0}(\check{x})=\tilde{f}_{0}\frac{\textstyle\sum_{l}\rho_{l}\Upsilon_{l}\big{(}\Phi_{x}^{\prime}(x_{0},\alpha_{l})\check{x}/|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|\big{)}}{\sum_{l}\rho_{l}},\\ &\Upsilon_{l}(r)=\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{u_{l}}^{u_{l}+r}\mathcal{F}^{-1}\bigg{(}\frac{1}{1+\kappa C_{1}|\lambda|^{3}}\bigg{)}(t)\text{d}t\text{d}u,\\ &\rho_{l}:=\frac{W^{2}(x_{0},y_{l})|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|}{|\Delta_{\Phi}(x_{0},\alpha_{l})|},\ u_{l}:=\frac{u}{|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|},\ C_{1}=\frac{1}{2\pi\textstyle\sum_{l}\rho_{l}}.\end{split}$
@@ 10. Numerical experiment
We choose $f$ to be the characteristic function of the disk centered at $x_{c}=(1,1)$ with radius $r=2$. Thus, $\mathcal{S}=\{x\in\mathbb{R}^{2}:\|x-x_{c}\|=r\}$. The center of a local region of interest (ROI) is $x_{0}=x_{c}+r\vec{\beta}_{0}$, where $\beta_{0}=0.17\pi$. The GRT integrates over circles $\mathcal{S}_{y}$, $y=(\alpha,\rho)$, with various radii $\rho>0$ and centers $R\vec{\alpha}$, $\alpha\in[0,2\pi)$, where $R=10$. Thus, $\rho=\Phi(x,\alpha)=\|x-R\vec{\alpha}\|$ and
(10.1) $\begin{split}&\Phi_{x}^{\prime}(x,\alpha)=\frac{x-R\vec{\alpha}}{\|x-R\vec{\alpha}\|}=:\vec{\Theta},\ \Phi_{\alpha}^{\prime}(x,\alpha)=R\vec{\Theta}\cdot\vec{\alpha}^{\perp},\\ &\Delta_{\Phi}(x,\alpha)=\det\begin{pmatrix}\vec{\Theta}&-\frac{R}{\|x-R\vec{\alpha}\|}\vec{\Theta}^{\perp}+c\vec{\Theta}\end{pmatrix}=-\frac{R}{\|x-R\vec{\alpha}\|}.\end{split}$
Here $\alpha^{\perp}=(-\sin\alpha,\cos\alpha)$, and $\Theta^{\perp}$ is defined similarly. The continuous data corresponds to $\alpha\in[0,2\pi)$ and $\rho\in(0,2R)$. The global reconstruction region is the square $\mathcal{U}_{b}:=\{(x_{1},x_{2})\in\mathbb{R}^{2}:|x_{1}|<R_{rec},\ |x_{2}|<R_{rec}\}$, $R_{rec}=3.7$. This corresponds to the set $\mathcal{U}_{b}$ used in (3.16). The discrete data are given at the points
@@
Clearly, for each $x\in\mathcal{S}$ there are two points, $y_{l}=(\alpha_{l},\rho_{l})$, $l=1,2$, such that $\mathcal{S}_{y_{l}}$ is tangent to $\mathcal{S}$ at $x$. To find $y_{1,2}$ for the selected $x_{0}$ we first solve $\|x_{c}+t\vec{\beta}_{0}\|=R$. This gives two values $t_{1},t_{2}$, $t_{2}<0<t_{1}$. Then $\alpha_{l}$ are determined from $x_{c}+t_{l}\vec{\beta}_{0}=R\vec{\alpha}_{l}$, and $\rho_{l}=\|x_{0}-R\vec{\alpha}_{l}\|$, $l=1,2$. In view of (10.1), in numerical computations we use $\|x_{0}-R\vec{\alpha}_{1}\|=t_{1}-r$ and $\|x_{0}-R\vec{\alpha}_{2}\|=-t_{2}+r$.
The regularization parameter is set at $\kappa=0.1$. In order to approximate the continuous GRT $\mathcal{R}$ and its adjoint $\mathcal{R}^{*}$, the reconstruction is performed on a dense $801\times 801$ grid covering the same square region $\mathcal{U}_{b}$, and the interpolated data $g_{\epsilon}$ are assumed to be given on a dense grid of the kind (10.2) with $N_{\alpha}^{\prime}=800$, $N_{\rho}^{\prime}=1201$. Prior to the reconstruction, the GRT data $g(y_{j})$ are interpolated from the grid (10.2) to the more dense grid using (3.13). For this we use the Keys interpolation kernel [34, 5]
@@
Minimization of the functional (3.16) is performed using gradient descent. Results of reconstructions are shown in Figures 2 and 3. Reconstruction of the entire region $\mathcal{U}_{b}$ is shown in Figure 2. A profile of the reconstruction through the center of the ball, $x_{c}$, is shown on the right. The location of the profile is shown in the left panel.
A small section of the reconstruction, which is located inside the square ROI shown in Figure 2, is extracted from the global reconstruction and shown in Figure 3. The right panel of Figure 3 shows the reconstructed (green) and predicted (blue) profiles. The latter is computed by using (10.1) in (3.20). The slight discrepancy in the interval $[1,2]$ is due to random numerical errors, which are clearly visible in the right panel of Figure 2. The plot in Figure 3 corresponds to the right jump in Figure 2. Overall, the match between the reconstruction and prediction is very accurate.
@@
It is easy to see that $\Psi:H_{c}^{1}(\mathcal{U}_{b})\to\mathbb{R}$ is strongly convex [7, Definition 1.1.48] and [41, Section 2.3] for each $\epsilon>0$. Strong convexity means that there exists $c>0$ such that
@@
for any $f_{1,2}\in H_{c}^{1}(\mathcal{U}_{b})$ and $\lambda\in(0,1)$. By an easy calculation,
(A.2) $\begin{split}\|\lambda\nabla f_{1}+(1-\lambda)\nabla f_{2}\|_{L^{2}(\mathcal{U})}^{2}\leq&\lambda\|\nabla f_{1}\|_{L^{2}(\mathcal{U})}^{2}+(1-\lambda)\|\nabla f_{2}\|_{L^{2}(\mathcal{U})}^{2}\\ &-2\lambda(1-\lambda)\|\nabla(f_{1}-f_{2})\|_{L^{2}(\mathcal{U})}^{2}\end{split}$
@@
(A.3) $\begin{split}\|\mathcal{R}(\lambda f_{1}&+(1-\lambda)f_{2})-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}\\ \leq&\lambda\|\mathcal{R}f_{1}-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}+(1-\lambda)\|\mathcal{R}f_{2}-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}\\ &-2\lambda(1-\lambda)\|\mathcal{R}(f_{1}-f_{2})\|_{L^{2}(\mathcal{V})}^{2}.\end{split}$
@@
(A.4) $\begin{split}\Psi(\lambda f_{1}&+(1-\lambda)f_{2})\leq\lambda\Psi(f_{1})+(1-\lambda)\Psi(f_{2})\\ &-c\lambda(1-\lambda)\big{(}\|\mathcal{R}(f_{1}-f_{2})\|_{L^{2}(\mathcal{V})}^{2}+\|\nabla(f_{2}-f_{1})\|_{L^{2}(\mathcal{U})}\big{)}.\end{split}$
@@
(A.6) $\|f\|_{H^{1}(\mathcal{U})}\leq c\big{(}\|f\|_{H^{-1/2}(\mathcal{U})}+\|\nabla f\|_{L^{2}(\mathcal{U})}\big{)},\ f\in C_{0}^{\infty}(\mathcal{U}_{b}).$
Applying (A.5) and (A.6) to (A.4) proves (A.1). Moreover, the functional is proper (its domain is not empty) and continuous $H_{c}^{1}(\mathcal{U}_{b})\to\mathbb{R}$. Thus the solution to (3.16) exists and is unique for each $\epsilon>0$ [7, Theorem 1.3.1], [41, Corollary 2.20].
@@
(B.1) $\begin{split}\partial_{p}H(x_{*}(y))=&\nu_{1*}(y)=\frac{|H_{x}^{\prime}(x_{*}(y))|}{|\Phi_{x}^{\prime}(x_{*}(y),\alpha)|}>0.\end{split}$
Another easy calculation, combined with (B.1) and (3.11), shows
(B.2) $\begin{split}\det M(y)&=-\big{(}[H_{xx}^{\prime\prime}(x_{*}(y))-\nu_{1*}(y)\Phi_{xx}^{\prime\prime}(x_{*}(y),\alpha)]e,e\big{)}|\Phi_{x}^{\prime}(x,\alpha)|^{2}\\ &=(\varkappa_{\mathcal{S}}(x)-\varkappa_{\mathcal{S}_{y}}(x)){|\Phi_{x}^{\prime}(x,\alpha)|^{2}}{|H_{x}^{\prime}(x)|},\end{split}$
where $e$ is a unit vector tangent to $\mathcal{S}$ at $x_{*}(y)$. Using assumption 3.3(2) proves the first statement.
Now we prove the second assertion in (4.5). Denote
(B.3) $e_{0}:=-\Phi_{x}^{\prime}(x_{*}(y),\alpha)/|\Phi_{x}^{\prime}(x_{*}(y),\alpha)|,\ V:=\begin{pmatrix}e_{0}&e\end{pmatrix}.$
Clearly, $V$ is a $2\times 2$ orthogonal matrix. By construction,
(B.4) $M_{1}:=\begin{pmatrix}1&0_{2}^{T}\\ 0_{2}&V^{T}\end{pmatrix}M\begin{pmatrix}1&0_{2}^{T}\\ 0_{2}&V\end{pmatrix}=\begin{pmatrix}0&|\Phi_{x}^{\prime}|&0\\ |\Phi_{x}^{\prime}|&*&*\\ 0&*&([H_{xx}^{\prime\prime}-\nu_{1*}\Phi_{xx}^{\prime\prime}]e,e)\end{pmatrix},$
$\text{sgn}M=\text{sgn}M_{1}$, and * denote various quantities, whose values are irrelevant. Next we use Haynsworth’s inertia additivity formula [18],
(B.5) $\begin{split}\text{sgn}M_{1}=&\text{sgn}\bigg{(}\big{(}\big{[}H_{xx}^{\prime\prime}-\nu_{1*}\Phi_{xx}^{\prime\prime}\big{]}e,e\big{)}-\begin{pmatrix}|\Phi_{x}^{\prime}|&0\end{pmatrix}\tilde{M}_{1}^{-1}\begin{pmatrix}|\Phi_{x}^{\prime}|\\ 0\end{pmatrix}\bigg{)}\\ &+\text{sgn}\tilde{M}_{1},\quad\tilde{M}_{1}:=\begin{pmatrix}0&|\Phi_{x}^{\prime}|\\ |\Phi_{x}^{\prime}|&*\end{pmatrix}.\end{split}$
An easy calculation shows that $\text{sgn}\tilde{M}_{1}=0$ and
(B.6) $\begin{split}\text{sgn}M_{1}=&\text{sgn}\big{(}\big{[}H_{xx}^{\prime\prime}-\nu_{1*}\Phi_{xx}^{\prime\prime}\big{]}e,e\big{)}=-1.\end{split}$
@@
Eq. (4.9) implies that $\mathcal{R}f\in I^{-3/2}(\mathcal{V},\Gamma)$ is a conormal (or, more generally, Lagrangian) distribution, see [22, Section 25.1]. By [22, Proposition 25.1.5], $\tilde{g}(\eta)=O(|\eta|^{-3/2})$, $|\eta|\to\infty$. Substituting this into (2.4) and using the equivalence of norms (2.3) and (2.4) gives $g\in C^{1/2}(\mathcal{V})$. As mentioned above, assumptions 3.1(1,3) imply that $\text{supp}\,g$ is compact.
The leading order term in (4.13) is obtained using [14, eq. 26, p. 360]:
(B.7) $\mathcal{F}((t-i0)^{a})=\frac{2\pi e(-a)}{\Gamma(-a)}\mu_{+}^{-(a+1)}.$
@@
Using the decay of the lower order terms in $\tilde{\upsilon}$ as $\lambda\to\infty$ (see $\tilde{R}$ in (4.9) and [22, eq. (25.1.4)]), gives $\Delta g\in C_{0}^{3/2}(\mathcal{V})$.
@@
It remains to prove the properties of the interpolated functions $g_{\epsilon}$ and $\Delta g_{\epsilon}$. Let $V\subset\mathbb{R}^{n}$ be any domain. Pick any $g\in C_{0}^{s}(V)$ and suppose first $10<s<$ 1. Let $\varphi\in C_{0}^{\lceil s\rceil}(\mathbb{R}^{n})$ be an interpolation kernel exact to degree $1$, and $g_{\epsilon}$ be the corresponding interpolated function. Denote $y_{h}:=y+h$, $\Delta g(y,h):=g(y_{h})-g(y)$ and similarly for $\Delta g_{\epsilon}(y,h)$. We have two equivalent representations
(B.8) $\begin{split}&\Delta g_{\epsilon}(y,h)-\Delta g(y,h)\\ &=\sum_{j}\varphi\bigg{(}\frac{y_{h}}{\epsilon}-j\bigg{)}\big{(}g(\epsilon j)-g(y_{h})\big{)}-\sum_{j}\varphi\bigg{(}\frac{y}{\epsilon}-j\bigg{)}\big{(}g(\epsilon j)-g(y)\big{)}\\ &=\sum_{j}\bigg{[}\varphi\bigg{(}\frac{y_{h}}{\epsilon}-j\bigg{)}-\varphi\bigg{(}\frac{y}{\epsilon}-j\bigg{)}\bigg{]}\big{(}g(\epsilon j)-g(y_{h})\big{)}\\ &\quad+\sum_{j}\varphi\bigg{(}\frac{y}{\epsilon}-j\bigg{)}\big{(}g(y_{h})-g(y)\big{)}.\end{split}$
We assume here for simplicity that $y_{j}=\epsilon j$, $j\in\mathbb{Z}^{2}$. When $h\geq\epsilon$, we use the second line above. Since $\varphi$ is compactly supported, the number of terms in each sum is finite. Moreover, $|\epsilon j-y_{h}|\leq c\epsilon$ in the first sum and $|\epsilon j-y|\leq c\epsilon$ in the second sum. Using that $g\in C_{0}^{s}(V)$, implies
(B.9) $|\Delta g_{\epsilon}(y,h)-\Delta g(y,h)|\leq ch^{s},\ h\geq\epsilon.$
When $0<h<\epsilon$, we use the second representation in (B.8). The kernel $\varphi$ is at least as smooth as $g$, so
(B.10) $\begin{split}\bigg{|}\frac{\varphi\big{(}(y_{h}/\epsilon)-j\big{)}-\varphi\big{(}(y/\epsilon)-j\big{)}}{(h/\epsilon)^{s}}\frac{g(\epsilon j)-g(y_{h})}{\epsilon^{s}}\bigg{|}<c.\end{split}$
The term on the last line in (B.8) is bounded by $ch^{s}$, and the assertion follows. The fact that $\|g_{\epsilon}\|_{L^{\infty}(V)}<c$ is obvious.
@@
Suppose now $g\in C_{0}^{s}(V)$ and $21<s<$ 2. As above, $\|g_{\epsilon}\|_{L^{\infty}(V)}<c$. By the exactness of $\varphi$ up to order one, differentiating (3.13) gives
@@
(B.11) $\begin{split}&g_{\epsilon}^{\prime}(y)=g^{\prime}(y)+\sum_{j}\varphi^{\prime}\bigg{(}\frac{y}{\epsilon}-j\bigg{)}\frac{g(\epsilon j)-(g(y)+(\epsilon j-y)\nabla g(y))}{\epsilon},\end{split}$

### v2

@@ 8.1. Reduction of the DTB function to an integral
(9.7) $\begin{split}&\Delta\mathcal{G}_{0}(\check{x})=\Delta f\frac{\textstyle\sum_{l}\nu_{l}\Upsilon_{l}\big{(}\text{d}_{x}\Phi(x_{0},\alpha_{l})\cdot\check{x}/|\text{d}_{x}\Phi(x_{0},\alpha_{l})|\big{)}}{\sum_{l}\nu_{l}},\\ &\Upsilon_{l}(r)=\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{u_{l}}^{u_{l}+r}\mathcal{F}^{-1}\bigg{(}\frac{1}{1+\kappa C_{1}|\lambda|^{3}}\bigg{)}(t)\text{d}t\text{d}u,\\ &\nu_{l}:=\frac{W^{2}(x_{0},y_{l})|\text{d}_{x}\Phi(x_{0},\alpha_{l})|}{|\Delta_{\Phi}(x_{0},\alpha_{l})|},\ u_{l}:=\frac{u}{|\text{d}_{x}\Phi(x_{0},\alpha_{l})|},\ C_{1}=\frac{1}{2\pi\textstyle\sum_{l}\nu_{l}}.\end{split}$
@@ 10. Numerical experiment
We choose $f$ to be the characteristic function of the disk centered at $x_{c}=(1,1)$ with radius $r=2$. Thus, $\mathcal{S}=\{x\in\mathbb{R}^{2}:|x-x_{c}|=r\}$. The center of a local region of interest (ROI), i.e. the point on the boundary $x_{0}\in\mathcal{S}$, is $x_{0}=x_{c}+r\vec{\beta}_{0}$, where $\beta_{0}=-0.17\pi$. The GRT integrates over circles $\mathcal{S}_{y}$, $y=(\alpha,\rho)$, with various radii $\rho>0$ and centers $R\vec{\alpha}$, $\alpha\in[0,2\pi)$, where $R=10$. The integration weight is $W\equiv 1$. Thus, $\rho=\Phi(x,\alpha)=|x-R\vec{\alpha}|$ and
(10.1) $\begin{split}&\text{d}_{x}\Phi(x,\alpha)=\frac{x-R\vec{\alpha}}{|x-R\vec{\alpha}|}=:\vec{\Theta},\ \Phi_{\alpha}^{\prime}(x,\alpha)=R\vec{\Theta}\cdot\vec{\alpha}^{\perp},\\ &\text{d}_{x}\Phi_{\alpha}^{\prime}(x,\alpha)=-\frac{R}{|x-R\vec{\alpha}|}\vec{\alpha}^{\perp}+c\vec{\Theta},\\ &\Delta_{\Phi}(x,\alpha)=\det\begin{pmatrix}\vec{\Theta}\\ -\frac{R}{|x-R\vec{\alpha}|}\vec{\alpha}^{\perp}+c\vec{\Theta}\end{pmatrix}=\frac{R\vec{\alpha}\cdot\vec{\Theta}}{|x-R\vec{\alpha}|}\not=0,\ |x|<R.\end{split}$
Here $\vec{\alpha}^{\perp}=(-\sin\alpha,\cos\alpha)$, and $\vec{\Theta}^{\perp}$ is defined similarly. The continuous data corresponds to $\alpha\in\mathcal{I}_{\alpha}:=[0,2\pi)\sim S^{1}$ and $\rho\in\mathcal{I}_{p}:=(0,2R)$. The image domain is $\mathcal{U}=\{|x|<R\}$. The support of the object is contained in the square reconstruction region $\mathcal{U}_{b}:=\{(x_{1},x_{2})\in\mathbb{R}^{2}:|x_{1}|<R_{rec},\ |x_{2}|<R_{rec}\}$, $R_{rec}=3.7$. This corresponds to the set $\mathcal{U}_{b}$ used in (3.17).
Further (cf. (3.22)),
(10.2) $\begin{split}&\nu_{l}=|\Delta_{\Phi}(x_{0},\alpha_{l})|^{-1},\ u_{l}=u,\ l=1,2.\end{split}$
This implies that the DTB function given by the fraction in (3.21) becomes
(10.3) $\frac{\textstyle\sum_{l}\nu_{l}\Upsilon_{l}(r)}{\textstyle\sum_{l}\nu_{l}}=\int_{\mathbb{R}}\frac{\textstyle\sum_{l}\nu_{l}h_{l}(u)}{\textstyle\sum_{l}\nu_{l}}\int_{u}^{u+r}R(t)\text{d}t\text{d}u.$
Let us now make sure that the circular GRT used in this section satisfies Assumptions 3.1 and 3.2. Clearly, Assumption 3.1 is satisfied. We next check Assumption 3.2. By the first two lines in (10.1), Assumption 3.2(1) is satisfied since $\vec{\alpha}\cdot\vec{\Theta}\not=0$ when $|x|<R$. Next, suppose $|x_{1}-R\vec{\alpha}|=|x_{2}-R\vec{\alpha}|$ for some $\vec{\alpha}$ and $x_{1}\not=x_{2}$, $|x_{1}|,|x_{2}|<R$. The first line in (10.1) implies $\Phi_{\alpha}^{\prime}(x_{1},\alpha)\not=\Phi_{\alpha}^{\prime}(x_{2},\alpha)$, therefore Assumption 3.2(2) is satisfied.
For any $(x_{0},\vec{\Theta}_{0})$, $|x_{0}|<R$, there are two circles $\mathcal{S}_{y}$ in the data, which contain $x_{0}$ and are orthogonal to $\vec{\Theta}_{0}$ at $x_{0}$. Thus, the conormal bundles of all the $\mathcal{S}_{y}$ cover $T^{*}\mathcal{U}_{b}$ and Assumption 3.2(3) is satisfied. Since $W\equiv 1$, Assumption 3.2(4) trivially holds.
By [6, Corollary 5], $\mathcal{R}$ is injective. As stated at the end of Section 3.1, together with the visibility of every singularity and the fact that $\mathcal{R}$ and $\mathcal{R}^{*}$ are FIOs of order $-1/2$, this ensures that Assumption 3.2(5) is satisfied.
The discrete data are given at the points
@@
As mentioned above, for each $x\in\mathcal{S}$ there are two points, $y_{l}=(\alpha_{l},\rho_{l})$, $l=1,2$, such that $\mathcal{S}_{y_{l}}$ is tangent to $\mathcal{S}$ at $x$. To find $y_{1,2}$ for the selected $x_{0}$ we first solve $\|x_{c}+t\vec{\beta}_{0}\|=R$. This gives two values $t_{1},t_{2}$, $t_{2}<0<t_{1}$. Then $\alpha_{l}$ are determined from $x_{c}+t_{l}\vec{\beta}_{0}=R\vec{\alpha}_{l}$, and $\rho_{l}=|x_{0}-R\vec{\alpha}_{l}|$, $l=1,2$. In view of (10.1), in numerical computations we use $|x_{0}-R\vec{\alpha}_{1}|=t_{1}-r$ and $|x_{0}-R\vec{\alpha}_{2}|=-t_{2}+r$. By a simple geometric argument (cf. (10.1)),
(10.5) $-R\vec{\alpha}_{l}\cdot\vec{\Theta}_{l}=(t_{1}-t_{2})/2,\ l=1,2.$
The regularization parameter is set at $\kappa=0.5$. In order to approximate the continuous GRT $\mathcal{R}$ and its adjoint $\mathcal{R}^{*}$, the reconstruction is performed on a dense $801\times 801$ grid covering the same square region $\mathcal{U}_{b}$, and the interpolated data $g_{\epsilon}$ are assumed to be given on a dense grid of the kind (10.4) with $N_{\alpha}^{\prime}=800$, $N_{\rho}^{\prime}=1201$. Prior to the reconstruction, the GRT data $g(y_{j})$ are interpolated from the grid (10.4) to the more dense grid using (3.14). For this we use the Keys interpolation kernel [37, 8]
@@
The minimization of the functional (3.17) is performed using gradient descent. The stopping criterion is met when the $L^{\infty}$-norm of the image update is less than $10^{-6}$ for three consecutive iterations. The algorithm converged in 214 iterations. The reconstruction results are shown in Figures 2 and 3. The reconstruction of the entire region $\mathcal{U}_{b}$ is shown in Figure 2. A profile of the reconstruction through the center of the ball, $x_{c}$, is shown on the right. The location of the profile is shown in the left panel.
A small section of the reconstruction, which is located inside the square ROI shown in Figure 2, is extracted from the global reconstruction and shown in Figure 3. The right panel of Figure 3 shows the reconstructed (green) and predicted (blue) profiles. The latter is computed by using (10.1) in (3.21), (3.22). The plot in Figure 3 corresponds to the right jump in Figure 2. Overall, the match between the reconstruction and prediction is very accurate.
@@
As the following shows, $\Psi:H_{0}^{1}(\mathcal{U}_{b})\to\mathbb{R}$ is strongly convex [9, Definition 1.1.48] and [45, Section 2.3] for each $\epsilon>0$. Indeed, strong convexity means that there exists $c>0$ such that
@@
for any $f_{1,2}\in H_{0}^{1}(\mathcal{U}_{b})$ and $\lambda\in(0,1)$. By an easy calculation,
(A.2) $\begin{split}\|\lambda\partial_{x}f_{1}+(1-\lambda)\partial_{x}f_{2}\|_{L^{2}(\mathcal{U})}^{2}\leq&\lambda\|\partial_{x}f_{1}\|_{L^{2}(\mathcal{U})}^{2}+(1-\lambda)\|\partial_{x}f_{2}\|_{L^{2}(\mathcal{U})}^{2}\\ &-\lambda(1-\lambda)\|\partial_{x}(f_{1}-f_{2})\|_{L^{2}(\mathcal{U})}^{2}\end{split}$
@@
(A.3) $\begin{split}\|\mathcal{R}(\lambda f_{1}&+(1-\lambda)f_{2})-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}\\ \leq&\lambda\|\mathcal{R}f_{1}-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}+(1-\lambda)\|\mathcal{R}f_{2}-g_{\epsilon}\|_{L^{2}(\mathcal{V})}^{2}\\ &-\lambda(1-\lambda)\|\mathcal{R}(f_{1}-f_{2})\|_{L^{2}(\mathcal{V})}^{2}.\end{split}$
@@
(A.4) $\begin{split}\Psi(\lambda f_{1}&+(1-\lambda)f_{2})\leq\lambda\Psi(f_{1})+(1-\lambda)\Psi(f_{2})\\ &-\lambda(1-\lambda)\big{(}\|\mathcal{R}(f_{1}-f_{2})\|_{L^{2}(\mathcal{V})}^{2}+\|\partial_{x}(f_{2}-f_{1})\|_{L^{2}(\mathcal{U})}\big{)}.\end{split}$
@@
(A.6) $\|f\|_{H^{1}(\mathcal{U})}\leq c\big{(}\|f\|_{H^{-1/2}(\mathcal{U})}+\|\partial_{x}f\|_{L^{2}(\mathcal{U})}\big{)},\ f\in C_{0}^{\infty}(\mathcal{U}_{b}).$
Applying (A.5) and (A.6) to (A.4) proves (A.1). Moreover, the functional is proper (its domain is not empty) and continuous $H_{0}^{1}(\mathcal{U}_{b})\to\mathbb{R}$. Thus the solution to (3.17) exists and is unique for each $\epsilon>0$ [9, Theorem 1.3.1], [45, Corollary 2.20].
@@
(B.1) $\begin{split}\partial_{p}H(x_{*}(y))=&\nu_{1*}(y)=\frac{|\text{d}_{x}H(x_{*}(y))|}{|\text{d}_{x}\Phi(x_{*}(y),\alpha)|}>0.\end{split}$
The last equality follows from (3.12) and the second equation in (4.4).
Let $e$ be a unit vector tangent to $\mathcal{S}$ at $x_{*}(y)$. Denote
(B.2) $e_{0}:=-\partial_{x}\Phi(x_{*}(y),\alpha)/|\partial_{x}\Phi(x_{*}(y),\alpha)|,\ V:=\begin{pmatrix}e_{0}&e\end{pmatrix}.$
Clearly, $V$ is a $2\times 2$ orthogonal matrix. Recall that $M$ is defined in (4.3). As is easily checked,
(B.3) $M_{1}:=\begin{pmatrix}1&0_{2}^{T}\\ 0_{2}&V^{T}\end{pmatrix}M\begin{pmatrix}1&0_{2}^{T}\\ 0_{2}&V\end{pmatrix}=\begin{pmatrix}0&|\text{d}_{x}\Phi|&0\\ |\text{d}_{x}\Phi|&a&b\\ 0&b&\text{d}_{x}^{2}[H-\nu_{1*}\Phi](e,e)\end{pmatrix},$
where $a,b$ are some quantities, whose values are irrelevant. Therefore, $\det M_{1}=\det M$ and $\text{sgn}M=\text{sgn}M_{1}$.
Using (B.1) and (3.11) gives
(B.4) $\begin{split}\det M(y)&=-\text{d}_{x}^{2}[H(x_{*}(y))-\nu_{1*}(y)\Phi(x_{*}(y),\alpha)](e,e)|\text{d}_{x}\Phi(x,\alpha)|^{2}\\ &=(\varkappa_{\mathcal{S}}(x)-\varkappa_{\mathcal{S}_{y}}(x)){|\text{d}_{x}\Phi(x,\alpha)|^{2}}{|\text{d}_{x}H(x)|}.\end{split}$
Assumption 3.3(2) completes the proof of the first statement.
To prove the second assertion in (4.5) we use Haynsworth’s inertia additivity formula [21],
(B.5) $\begin{split}\text{sgn}M_{1}=&\text{sgn}\bigg{(}\text{d}_{x}^{2}[H-\nu_{1*}\Phi](e,e)-{\color[rgb]{0,0,0}\begin{pmatrix}0&b\end{pmatrix}}\tilde{M}_{1}^{-1}{\color[rgb]{0,0,0}\begin{pmatrix}0\\ b\end{pmatrix}}\bigg{)}\\ &+\text{sgn}\tilde{M}_{1},\quad\tilde{M}_{1}:=\begin{pmatrix}0&|\text{d}_{x}\Phi|\\ |\text{d}_{x}\Phi|&a\end{pmatrix}.\end{split}$
An easy calculation shows that
(B.6) $\text{sgn}\tilde{M}_{1}=0,\quad\begin{pmatrix}0&b\end{pmatrix}\tilde{M}_{1}^{-1}\begin{pmatrix}0\\ b\end{pmatrix}=b^{2}\big{(}\tilde{M}_{1}^{-1}\big{)}_{22}=0.$
Therefore, arguing as in (B.4) and using Assumption 3.3(2), we conclude that
(B.7) $\begin{split}\text{sgn}M_{1}=&\text{sgn}\big{(}\text{d}_{x}^{2}[H-\nu_{1*}\Phi](e,e)\big{)}=-1.\end{split}$
@@
Eqs. (4.9), (4.10) imply that $\hat{f}=\mathcal{R}f\in I^{-3/2}(\mathcal{V},\Gamma)$ is a conormal (or, more generally, Lagrangian) distribution, see [26, Section 25.1]. By [16, Lemma 2.1 and its proof] with $\mu=-3/2$ and $k=1$, $g=\chi_{V}\hat{f}\in C_{0}^{1/2}(V)$. The leading order term in (4.13) is obtained using [17, eq. 26, p. 360]:
(B.8) $\mathcal{F}((t-i0)^{-\sigma})=\frac{2\pi e(\sigma)}{\Gamma(\sigma)}\mu_{+}^{\sigma-1},\ \sigma\in\mathbb{R},$
where $\mu$ is the Fourier transform variable.
@@
Using the decay of the lower order terms in $\tilde{\upsilon}$ as $\lambda\to\infty$ (namely, $\tilde{R}$ in (4.10)), gives $\Delta g\in C_{0}^{3/2}(V)$. Here we apply [16, Lemma 2.1 and its proof] with $\mu=-3/2$ and $k=1$ to $\partial_{p}\Delta g$ and $\partial_{\alpha}\Delta g$ and then use the first line in (2.3) with $s=3/2$ and $k=1$.
@@
It remains to prove the properties of the interpolated functions $g_{\epsilon}$ and $\Delta g_{\epsilon}$. Pick any $g\in C_{0}^{s}(V)$. Let $\varphi\in C_{0}^{\lceil s\rceil}(\mathbb{R}^{2})$ be an interpolation kernel exact to degree $1$, and $g_{\epsilon}$ be the corresponding interpolated function. Denote $y_{h}:=y+h$. We assume here for simplicity that $y_{j}=\epsilon j$, $j\in\mathbb{Z}^{2}$.
Suppose first $10<s<$ 1. If $|h|>\epsilon$, it suffices to prove that $|g_{\epsilon}(y)-g(y)|\leq c\epsilon^{s}$. Then $|g_{\epsilon}(y_{h})-g_{\epsilon}(y)|\leq c|h|^{s}$ follows from the triangle inequality and the assumption $g\in C_{0}^{s}(V)$. By the exactness of $\varphi$,
(B.9) $\begin{split}g_{\epsilon}(y)-g(y)=\sum_{j}\varphi\bigg{(}\frac{y-\epsilon j}{\epsilon}\bigg{)}\big{(}g(\epsilon j)-g(y)\big{)}.\end{split}$
Since $\varphi$ is compactly supported, $|\epsilon j-y|\leq c\epsilon$ and the number of terms in the sum is finite, and the desired assertion follows.
If $|h|<\epsilon$, we have
(B.10) $\begin{split}&g_{\epsilon}(y_{h})-g_{\epsilon}(y)=\sum_{j}\bigg{[}\varphi\bigg{(}\frac{y_{h}-\epsilon j}{\epsilon}\bigg{)}-\varphi\bigg{(}\frac{y-\epsilon j}{\epsilon}\bigg{)}\bigg{]}\big{(}g(\epsilon j)-g(y)\big{)}.\end{split}$
By assumption, $\varphi$ is at least as smooth as $g$, so
(B.11) $\begin{split}\bigg{|}\frac{\varphi\big{(}(y_{h}-\epsilon j)/\epsilon\big{)}-\varphi\big{(}(y-\epsilon j)/\epsilon\big{)}}{(|h|/\epsilon)^{s}}\frac{g(\epsilon j)-g(y)}{\epsilon^{s}}\bigg{|}<c.\end{split}$
Hence $|g_{\epsilon}(y_{h})-g_{\epsilon}(y)|\leq c|h|^{s}$. The fact that $\|g_{\epsilon}\|_{L^{\infty}(V)}<c$ is obvious.
@@
Suppose now $g\in C_{0}^{s}(V)$, $21<s<$ 2, i.e., we study $\Delta g$ of Lemma 4.2. As above, $\|g_{\epsilon}\|_{L^{\infty}(V)}<c$. By the exactness of $\varphi$ up to order one, differentiation of (3.14) gives
@@
(B.12) $\begin{split}&g_{\epsilon}^{\prime}(y)=g^{\prime}(y)+\sum_{j}\varphi^{\prime}\bigg{(}\frac{y-\epsilon j}{\epsilon}\bigg{)}\frac{g(\epsilon j)-\big{[}g(y)+\text{d}_{y}g(y)\cdot(\epsilon j-y)\big{]}}{\epsilon},\end{split}$